In each of the following cases, find the unknown variable. Ignore taxes. (Do not round intermediate calculations and round your answers to the nearest whole number, e.g., 32.)

Accounting
Break-Even
Unit Price Unit Variable Cost Fixed Costs Depreciation
18,000     $ 42 $ 22 $ 250,000 $
128,000     42 4,200,000 920,000
7,600     75 300,000 80,000
Answer #1

AccountingBreak-Even=(Fixed cost+Depreciation)/(Unit price-Unit Variable cost)

1.

18000=(250,000+Depreciation)/(42-22)

18000*20=250,000+Depreciation

Depreciation=360,000-250,000

=$110,000.

2.

128000=(4,200,000+920000)/(Unit price-42)

128000*(Unit price-42)=5,120,000

Unit price-42=(5,120,000/128000)

Unit price=40+42

=$82

3.

7600=(300,000+80,000)/(75-Variable cost)

7600*(75-Variable cost)=380,000

75-Variable cost=(380,000/7600)

Variable cost=75-50

=$25